OpenStudy (anonymous):

How much money should you invest now to have $7000 in 12 years if you invest the money at a rate of 9.5%, compounded semiannually?

OpenStudy (anonymous):

A=P(1+r/n)^(nt) P=A/[(1+r/n)^(nt)} P=7000/[1+.095/2)^(12*2)]=$2,298.27